OpenStudy (purplemexican):

Aiko is finding the sum (4 + 5i) + (–3 + 7i). She rewrites the sum as (–3 + 7)i + (4 + 5)i. Which statement explains the property of addition that she made an error using? Aiko incorrectly used the commutative property by changing the order of the two complex numbers. Aiko incorrectly used the identity property by combining the real number and the coefficient of the imaginary part. Aiko incorrectly used the distributive property by combining the real number and the coefficient of the imaginary p

OpenStudy (purplemexican):

ok could you explain how i do this so i can figure it out on my own?

OpenStudy (michele_laino):

we have this: \[\Large \left( { - 3 + 7} \right)i + \left( {4 + 5} \right)i = - 3i + 7i + 4i + 5i = 13i\] which is not right since it is a purely imaginary number, whereas the sum of your originals number is: \[\Large b\left( {4{\text{ }} + {\text{ }}5i} \right) + \left( {--3{\text{ }} + {\text{ }}7i} \right) = \left( {4 - 3} \right) + \left( {5 + 7} \right)i\] which is not a purely imaginary number

OpenStudy (michele_laino):

oops.. \[\Large \left( {4 + 5i} \right) + \left( { - 3 + 7i} \right) = \left( {4 - 3} \right) + \left( {5 + 7} \right)i\]

OpenStudy (purplemexican):

thank you for the help but its to late now

OpenStudy (michele_laino):

ok! thanks! :)